OpenStudy (y2o2):

prove that :

OpenStudy (y2o2):

\[\large {^{2n} C _n} = {(^n C _0)}^2 +{(^n C _1)}^2 +{(^n C _2)}^2 +...+{(^n C _n)}^2 \]

OpenStudy (kinggeorge):

So in summation notation, show that\[\binom{2n}{n}=\sum_{i=0}^n \binom{n}{i}^2\]

OpenStudy (shubhamsrg):

you can use induction..

OpenStudy (zarkon):

Use Vandermonde's identity

OpenStudy (mimi_x3):

well have you tried it?

OpenStudy (mimi_x3):

since it was from 2 months ago..

OpenStudy (y2o2):

Yes , and it came up with nothing

OpenStudy (anonymous):

no it works.

OpenStudy (mimi_x3):

\[LHS: \binom{2n}{n} x^{n} \] \[RHS\left[\binom{n}{0}\binom{n}{n}+\binom{n}{1}\binom{n}{n-1}+\binom{n}{2}\binom{n}{n-2}+....+\binom{n}{n-2}\binom{n}{2}+\binom{n}{n-1}\binom{n}{1}+\binom{n}{n}\binom{n}{0}\right] *x^n\] \[By ~Symmetry: \binom{n}{n} = \binom{n}{0} ; \binom{n}{n-1} = \binom{n}{1} \] \[hence~ RHS = \left[\binom{n}{0}^{2}+\binom{n}{1}^{2}+\binom{n}{2}^{2}+...+\binom{n}{n}^{2}\right] x^n\] \[hence, coeeficients~ of~ x^n~ are~ the ~same \] \[\therefore, \binom{n}{0}^{2}+\binom{n}{1}^{2}+\binom{n}{2}^{2}+...+\binom{n}{n}^{2} = \binom{2n}{n} \]

OpenStudy (mimi_x3):

whats wrong with the method i used?

OpenStudy (anonymous):

nothing. I said its easier to use vandermonde's identity.

OpenStudy (mimi_x3):

okay; well i dont know how to use that identity lol

OpenStudy (anonymous):

just plug in n every where except at k(its a summation variable).

OpenStudy (y2o2):

thank you, i got it now :)

OpenStudy (mimi_x3):

You're welcome (:

OpenStudy (mimi_x3):

Sorry I made a typo at the end: \[\therefore, \binom{n}{0}^{2}+\binom{n}{1}^{2}+\binom{n}{2}^{2}+...+\binom{n}{n}^{2} = \binom{2n}{n} ^2\]

OpenStudy (y2o2):

yeah , i realized that :)